OpenStudy (anonymous):

find the LCM x^2+2x,x^2-2x

OpenStudy (mathstudent55):

Factor each binomial. Then you need one factor of each of all the factors for the LCM

OpenStudy (anonymous):

Is the answer x^2 (x+2)(x-2)

OpenStudy (anonymous):

Factors x^2+2x : x (x+2) x^2-2x :x (x-2) LCM is product of one factor from each column = x(x+2)(x-2)

OpenStudy (anonymous):

Thank you. I thought it I had to put x^2 (x+2)(x-2)

OpenStudy (mathstudent55):

x^2 + 2x = x(x + 2) x^2 - 2x = x(x - 2) For LCM, you need common and non-common factors with the largest exponent. x is a factor common to both, so take x x + 2 is not common to both, so take it also x - 2 is not common to both, so take it also LCM = x(x + 2)(x -2)